Higher is a producer of kites and wind socks. Relevant data on the bottleneck operation in the shop for the upcoming fiscal year are given in the following table:

Item Kites Wind Socks
demand forecast 30,000 units/yr 15,000 units/yr
lot size 20 units 60 units
standard processing time 0.4 hours/unit 0.5 hours/unit
standard setup time 3 hours/lot 4 hours/lot

The shop works two shifts per day, 8 hours per shift, 180 days per year. Currently, the company operates 4 machines, and desires a 25 percent capacity cushion. How many machines should be purchased to meet the upcoming year's demand without resorting to any short-term capacity solutions?
Answer #1

Answer:

  • Total machining hours per machine per year= no. of shifts per day* duration of shift * no. of days per year
  • Total machining hours per year= A = 2*8*180 = 2880 hours

Kites:

  • Total no. of lots = B = Demand forecast / lot size = 30000 /20 = 1500 lots
  • Total time required to manufacture demand = Demand *per unit time
  • Total time required to manufacture demand = C = 30000*0.4 = 12000 hours
  • Total setup time = D =No. of lots *set up time = 1500*3 = 4500 hours
  • Total time required for Kites = E = C+D = 12000+4500= 16500 Hours

Wind Socks:

  • Total no. of lots = F = Demand forecast / lot size = 15000 /60 = 250 lots
  • Total time required to manufacture demand = Demand *per unit time
  • Total time required to manufacture demand = G = 15000*0.5 = 7500 hours
  • Total setup time = H =No. of lots *set up time = 250*4 = 1000 hours
  • Total time required for Kites = I = G+H = 7500+1000= 8500 Hours
  • Total Machining hours required =J = E+I = 16500 + 8500 =25000 hours
  • Total Machining hours required considering capacity cushion of 25% = K = J/0.75 = 33333.33 hours
  • No. of machines required to meet "K" = K/ A = 11.57machines
  • But we have 4 machines , so we will require 7.57 more machines.
  • Hence the no. of machines to be purchased=7.57= 8 (Answer)